# FPGA-based Emulation and Device-Side Management for CXL-based Memory Tiering Systems

### v2

@@ 4.4. Migration Unit
4.5. Software Interface
HeteroMem can manage the memory tiering system transparently without CPU intervention. To better understand the behavior of HeteroMem, we have implemented a series of profiling registers, each indicating the number of specific events. Additionally, HeteroMem’s behavior can be tuned through several parameters, which we have also implemented as configuration registers. We expose these configuration and profiling registers in the PCIe BAR space of the CXL memory device. The host CPU can read or write these registers through MMIO. In the OS kernel space, we have implemented a device driver that exposes the MMIO address of these registers to user space as files under the $/sys/kernel/mm/heteromem$ directory. Users can read or write these files to configure HeteroMem and obtain profiling information. For example, users can read $/sys/kernel/mm/heteromem/migrated\_pages\_cnt$ to get the number of pages migrated by HeteroMem.
5. Implementation
Table 2. Evaluation System Configuration
@@ 5. Implementation
Table 3. Parameters Configuration
Parameter Value Description
Sketch Width(W) 512k Width of the sketch in hotness profiling logic.
Sketch Depth(D) 2 Depth of the sketch in hotness profiling logic.
sketch_reset_period $50$-$5000ms$ The reset period of the sketch in HeteroMem.
bitmap_reset_period $1$-$5s$ The reset period of the bitmap in HeteroMem.
migrate_page_limit $256MB/s$ The limit of migration speed for all memory tiering systems.
pebs_sampling_rate $397$-$19997$ The sampling rate of PEBS
pte_sampling_rate $1$-$3s$ PTE-sampling rate of TPP&AutoNUMA
page_scanning_rate $5$-$10s$ Page table scanning rate of PTE-scan
Figure 8. FPGA-based Prototyping System
@@
We implement HeteroMem based on the Linux kernel v6.3. The CXL memory on the FPGA is exposed to the software as a CPU-less NUMA node. To compare the performance of HeteroMem with other baseline memory tiering systems, we modified the NUMA nodes enumeration code and split the CXL memory NUMA node into two fake NUMA nodes, corresponding to the fast memory and slow memory emulated by HeteroBox. We use Intel® FPGA CXL IP (Memory Expander Type 3) and add about 7000 lines of Verilog RTL custom logic code to build HeteroBox and HeteroMem, along with approximately 1000 lines of C++ simulation environment code and about 1000 lines of modifications in the Linux kernel for HeteroBox and HeteroMem driver and fake NUMA nodes logic. We will open-source all the hardware and software codes after publication.
Figure 9. Validation of HeteroBox Emulation Platform
@@
6. Evaluation
6.1. Experimental Setup
@@
Baselines: We compare the performance of our HeteroMem design with five baseline memory tiering systems: PTE-Scan (Park, 2021), PEBS (peb, [n. d.]), Auto NUMA (Corbet, 2012), TPP (Maruf et al., 2023), and NeoMem (Zhou et al., 2024). PTE-Scan utilizes the access bit in the page table entry for hotness profiling. The access bit is set by the MMU when a page is accessed, and PTE-Scan periodically scans the page table, clearing the access bit and detecting a page as hot if its access bit is set. PTE-Scan migrates hot pages to fast memory as soon as they are detected. PEBS uses hardware counters to profile page hotness. It is configured to sample the $MEM\_LOAD\_L3\_MISS\_RETIRED$ event. When an LLC miss occurs, the hardware counters increment by one. When the counter reaches a given threshold, it overflows and records the address of the current memory access. The PEBS memory tiering system detects a page as hot and migrates it to fast memory when its address is recorded. TPP and AutoNUMA utilize hint-fault monitoring to detect hot data in slow memory. NeoMem is a recently proposed memory tiering system that achieves state-of-the-art performance. It enhances the CXL memory device with a hot page profiling structure, which identifies and reports hot pages to the host CPU for migration to faster memory. Table 3 presents the parameter configurations for these memory tiering systems.
@@
Benchmarks: We select eight benchmarks for evaluation: GUPS (gup, 2021), a microbenchmark that performs parallel read-modify-write operations in a uniform or skewed random pattern within its working set; Silo (Tu et al., 2013), an in-memory database engine; Btree, an in-memory index lookup; XSBench (Tramm et al., 2014), an HPC workload; Betweenness Centrality (BC) and PageRank (PR)(Beamer et al., 2017), two graph processing workloads; and 603.bwaves and 654.roms, two benchmarks from SPEC 2017. These benchmarks are widely used to evaluate memory tiering systems in previous works(Raybuck et al., 2021; Lee et al., 2023; Zhou et al., 2024). The Resident Set Size (RSS) of these benchmarks ranges from 4.4GB to 11.1GB.
6.2. Validation of HeteroBox
@@
We use the Intel MLC tool (mlc, [n. d.]) to evaluate the latency and bandwidth attribute of the HeteroBox emulation platform. We first focus on the case of a single region. As illustrated in Figure 9-(a), we configure the HeteroBox to contain one region and set the latency register value from 0 to 256. Using the Intel MLC, we then measure the region’s latency. Changing the bandwidth configuration values from 128 to 32, we plot four curves of relationship between the value of latency register and the latency attribute of this region. Additionally, we include a curve called “theoretical latency”, which incorporates the latency from both the CXL link and the CXL IP on the FPGA. The theoretical latency also includes the latency introduced by the HeteroBox emulation logic. Due to the latency introduced by the underlying memory controller and DDR4 DRAM, the region’s latency is the same when the latency register is set to 0 or 32. All four curves corresponding to different bandwidth configuration values overlap, showing consistent latency performance when compared to theoretical latency. These results demonstrate two key findings: (1) The HeteroBox emulation logic can precisely adjust latency attribute of a region. (2) Modifying the bandwidth configuration register value does not affect the latency attribute of the region.
@@
Figure 10. Analysis of Remapping Overhead.
Figure 11. Performance Comparison with Different Fast Memory Size Configuration.
@@
6.3. Remapping Overhead
Then we discuss the overhead introduced by HeteroMem. The Profiling Unit of HeteroMem is off the critical path, so it does not introduce extra latency. The Migration Unit only operates when a migration transaction occurs. Since we limit the number of pages migrated within a given period (cycles consumed on migration are limited to below 5%), the extra latency introduced by Migration Unit is minimal. However, Remapping Unit is always on the critical path and incurs non-negligible extra latency and bandwidth overhead when the remapping cache misses. As a result, the overhead introduced by the Remapping Unit should be carefully considered.
As shown in Figure 10-(a), we disable migration in HeteroMem and test the performance slowdown with the Remapping Unit enabled, varying the remapping cache size from 128KB to 2MB. Figure 10-(b) displays the remapping cache hit rate for different cache sizes. For all benchmarks, both the performance and the remapping cache hit rate improve as the remapping cache size increases. The GUPS benchmark shows a low remapping cache hit rate with smaller cache sizes, causing significant overhead due to its random memory access pattern. This results in a low cache hit rate when the cache cannot cover the accessed memory range. However, as the remapping cache size increases and covers the memory range accessed by GUPS, the hit rate improves significantly, reaching nearly 100% at 2MB, reducing the Remapping Unit’s overhead. For BC, PR and XSBench benchmarks, the remapping cache hit rate also approaches 100%, resulting in minimal overhead with a 2MB cache. Conversely, the Btree and Silo benchmarks exhibit lower locality, leading to low remapping cache hit rates even with a 2MB cache, and show performance slowdowns of 11.8% and 11.0%, respectively. Results from subsequent experiments indicate that the performance degradation in these two applications caused by Remapping Unit can be offset by HeteroMem’s superior memory tiering solution.
6.4. Main Results
@@
Figure 12. Sensitivity Analysis and Profiling Result of HeteroMem
Figure 13. Analysis of Memory Access Distribution.
As shown in Figure 11, HeteroMem consistently maintains superior geomean performance compared to baseline systems across various fast memory sizes. All performance statistics are normalized to NeoMem. In the 4GB fast memory configuration, HeteroMem’s geomean performance is 5.1% to 16.2% higher than baseline systems. For BC benchmark, HeteroMem achieves 32.7% higher performance than NeoMem and PTE-Scan, and 44.6% higher than PEBS. Additionally, the performance improvement of HeteroMem increases with larger fast memory sizes, primarily due to the enhanced effectiveness of HeteroMem’s accurate cold page profiling mechanism. The performance improvement of HeteroMem stems from its accurate hardware-based profiling method and efficient hardware-based data movement mechanism, as described in Section 4.
@@
In Figure 13-(b), we plot the memory access distribution for the BC benchmark. We implement a 32-bit hardware counter for each 2MB physical page in the CXL extended memory and record the memory access count to each 2MB page every second. The results show that HeteroMem quickly migrates all hot data to fast memory, whereas the baseline systems leave some hot data in slow memory, demonstrating the effectiveness of HeteroMem’s memory management strategy.
7. Related Work
7.1. Software-based memory tiering system
@@
Many previous works using software-based methods have tried to optimize the performance of memory tiering systems in various aspects, including memory access profiling (Agarwal and Wenisch, 2017; Choi et al., 2021; Lee and Eom, 2022; Park, 2021), page migration (Kommareddy et al., 2019; Ryoo et al., 2018; Yan et al., 2019), page classification (Koh et al., 2019; Maruf et al., 2022), and huge page strategies (Heo et al., 2022; Lee et al., 2023). TPP (Maruf et al., 2023) proposes to promote a page after it is scanned with the access bit set twice. MEMTIS (Lee et al., 2023) uses histogram-based hot set classification to achieve more stable hot page detection. NOMAD (Xiang et al., 2024) employs transactional page migration to mitigate the overhead of page migration. Additionally, there are works that utilize application-level information to manage data placement at the object level (mem, 2021; pmd, 2017; Hildebrand et al., 2020; Li and Wu, 2022; Ren et al., 2021; Wang et al., 2019; Wei et al., 2015). However, although these works strive to optimize the performance of memory tiering systems, their performance is limited by the inherent inaccuracy of software profiling and the high overhead of data migration using the CPU, as described in Section 2. Our HeteroMem design employs a hardware-based high-resolution profiling scheme for hot and cold page detection. Meanwhile, HeteroMem uses hardware-based data movement, which achieves high bandwidth and low overhead without wasting CPU cycles and CXL link bandwidth.
7.2. Hardware-based memory tiering system
@@
There are also many works that try to optimize memory tiering systems using hardware-based methods. They manage fast memory as a memory cache (opt, [n. d.]) or use fast memory in flat memory mode (Chou et al., 2014; Ryoo et al., 2017; Kotra et al., 2018; Dong et al., 2010b; Prodromou et al., 2017; Li et al., 2023b). Intel Optane persistent memory provides memory mode (opt, [n. d.]), allowing users to use DRAM as a cache for non-volatile memory. CAMEO (Chou et al., 2014) combines memory cache mode and flat address mode. It manages fast memory in a cache-like manner, constrains data movement within associated sets, and moves data at cacheline granularity. MemPod (Prodromou et al., 2017) groups HBM and DRAM connected to CPU into clusters called Pods and manages data placement within each Pod. However, these hardware-based methods, proposed before the CXL era, involve intrusive modifications to CPU, making them inflexible and unable to adapt to dynamic and diverse CXL-based device-side heterogeneous memory systems. HeteroMem, on the other hand, places all its hardware logic on the device side and can be runtime configured to adapt to various heterogeneous memory system configurations, offering a one-for-all solution.
@@
8. Discussion & Future Work
Support for CXL memory pooling and sharing: Memory pooling and sharing are promising use cases of CXL (Ha et al., 2023), allowing memory to be shared between processors (consortium, [n. d.]). Memory in a memory pool can be dynamically allocated and deallocated, leading to more efficient use of memory resources. Memory sharing, on the other hand, enables each processor to read from or write to the shared memory, facilitating easy data sharing and communication between processors. The data migration process of HeteroMem is completely transparent to CPU. As a result, HeteroMem serves as a plug-in solution for both memory pooling and sharing scenarios.
Hardware overhead and Scalability: The hardware overhead of HeteroMem remains stable as the CXL-extended memory system scales. Firstly, HeteroMem’s hardware overhead increases minimally with the number of CXL devices. HeteroMem can be integrated into a CXL switch, and the number of CXL switches is significantly smaller than the number of CXL devices. Furthermore, not every CXL switch needs to be equipped with HeteroMem hardware. HeteroMem can manage the entire memory space it can access, so it only needs to be implemented in the CXL switch connected to the host CPU. Secondly, as the memory space scales, the increase in HeteroMem’s hardware overhead remains minimal. The hot page profiling component of HeteroMem utilizes Count-Min Sketch to detect hot pages, with an error bound related to the memory access sequence length rather than the memory space size. The cold page profiling component, namely the ping-pong bitmap, uses two bits of SRAM for every page. This structure only needs to be allocated for fast memory regions with demotion targets, which typically have a small capacity. For the Remapping Unit, the page granularity can be increased as memory space expands, such as using a mix of 4KB and 2MB pages, or exclusively 2MB pages. For example, a 4MB Remapping Cache (potentially larger in ASIC designs) with 2MB page granularity has a reach of 1TB memory (assuming 8 bytes per entry). This reach can be further extended by using larger page granularity (e.g., 1GB pages).

### v3

@@ IV-D Migration Unit
V Implementation
TABLE II: Evaluation System Configuration
@@ V Implementation
Figure 8: FPGA-based Prototyping System
@@
We implement HeteroMem based on the Linux kernel v6.3. The CXL memory on the FPGA is exposed to the software as a CPU-less NUMA node. To compare the performance of HeteroMem with other baseline memory tiering systems, we modified the NUMA nodes enumeration code and split the CXL memory NUMA node into two fake NUMA nodes, corresponding to the fast memory and slow memory emulated by HeteroBox. We use Intel® FPGA CXL IP (Memory Expander Type 3) and add about 7000 lines of Verilog RTL custom logic code to build HeteroBox and HeteroMem, along with approximately 1000 lines of C++ simulation environment code and about 1000 lines of modifications in the Linux kernel for HeteroBox and HeteroMem driver and fake NUMA nodes logic. The HeteroBox is publicly available in the provided repository111 https://github.com/memory-of-star/HeteroBox.
Figure 9: Validation of HeteroBox Emulation Platform
Figure 10: Performance Comparison with Different Fast Memory Size Configuration.
@@
VI Evaluation
VI-A Experimental Setup
@@
Baselines: We compare the performance of our HeteroMem design with five baseline memory tiering systems: PTE-Scan [41], PEBS [3], Auto NUMA [14], TPP [40], and NeoMem [61]. PTE-Scan utilizes the access bit in the page table entry for hotness profiling. The access bit is set by the MMU when a page is accessed, and PTE-Scan periodically scans the page table, clearing the access bit and detecting a page as hot if its access bit is set. PTE-Scan migrates hot pages to fast memory as soon as they are detected. PEBS uses hardware counters to profile page hotness. It is configured to sample the MEM_LOAD_L3_MISS_RETIREDMEM\_LOAD\_L3\_MISS\_RETIRED event. When an LLC miss occurs, the hardware counters increment by one. When the counter reaches a given threshold, it overflows and records the address of the current memory access. The PEBS memory tiering system detects a page as hot and migrates it to fast memory when its address is recorded. TPP and AutoNUMA utilize hint-fault monitoring to detect hot data in slow memory. NeoMem is a recently proposed memory tiering system that achieves state-of-the-art performance. It enhances the CXL memory device with a hot page profiling structure, which identifies and reports hot pages to the host CPU for migration to faster memory.
@@
Benchmarks: We select eight benchmarks for evaluation: GUPS [5], a microbenchmark that performs parallel read-modify-write operations in a uniform or skewed random pattern within its working set; Silo [54], an in-memory database engine; Btree, an in-memory index lookup; XSBench [53], an HPC workload; Betweenness Centrality (BC) and PageRank (PR)[8], two graph processing workloads; and 603.bwaves and 654.roms, two benchmarks from SPEC 2017. These benchmarks are widely used to evaluate memory tiering systems in previous works[43, 34, 61]. The Resident Set Size (RSS) of these benchmarks ranges from 4.4GB to 11.1GB.
VI-B Validation of HeteroBox
@@
We use the Intel MLC tool [2] to evaluate the latency and bandwidth attribute of the HeteroBox emulation platform. We first focus on the case of a single region. As illustrated in Figure 9-(a), we configure the HeteroBox to contain one region and set the latency register value from 0 to 256. Using the Intel MLC, we then measure the region’s latency. Changing the bandwidth configuration values from 128 to 32, we plot four curves of relationship between the value of latency register and the latency attribute of this region. Additionally, we include a curve called “theoretical latency”, which incorporates the latency from both the CXL link and the CXL IP on the FPGA. The theoretical latency also includes the latency introduced by the HeteroBox emulation logic. Due to the latency introduced by the underlying memory controller and DDR4 DRAM, the region’s latency is the same when the latency register is set to 0 or 32. All four curves corresponding to different bandwidth configuration values overlap, showing consistent latency performance when compared to theoretical latency. These results demonstrate two key findings: (1) The HeteroBox emulation logic can precisely adjust latency attribute of a region. (2) Modifying the bandwidth configuration register value does not affect the latency attribute of the region.
@@
Figure 11: Sensitivity Analysis and Profiling Result of HeteroMem
Figure 12: Analysis of Memory Access Distribution.
@@
VI-C Main Results
@@
As shown in Figure 10, HeteroMem consistently maintains superior geomean performance compared to baseline systems across various fast memory sizes. All performance statistics are normalized to NeoMem. In the 4GB fast memory configuration, HeteroMem’s geomean performance is 5.1% to 16.2% higher than baseline systems. For BC benchmark, HeteroMem achieves 32.7% higher performance than NeoMem and PTE-Scan, and 44.6% higher than PEBS. Additionally, the performance improvement of HeteroMem increases with larger fast memory sizes, primarily due to the enhanced effectiveness of HeteroMem’s accurate cold page profiling mechanism. The performance improvement of HeteroMem stems from its accurate hardware-based profiling method and efficient hardware-based data movement mechanism, as described in Section IV.
@@
In Figure 12-(b), we plot the memory access distribution for the BC benchmark. We implement a 32-bit hardware counter for each 2MB physical page in the CXL extended memory and record the memory access count to each 2MB page every second. The results show that, compared to the baseline systems, HeteroMem quickly and accurately migrates all hot data to fast memory, demonstrating the effectiveness of its memory management strategy.
VII Related Work
VII-A Software-based memory tiering system
@@
Many previous works using software-based methods have tried to optimize the performance of memory tiering systems in various aspects, including memory access profiling [7, 11, 33, 41], page migration [30, 46, 58], page classification [29, 39], and huge page strategies [24, 34]. TPP [40] proposes to promote a page after it is scanned with the access bit set twice. MEMTIS [34] uses histogram-based hot set classification to achieve more stable hot page detection. NOMAD [57] employs transactional page migration to mitigate the overhead of page migration. Additionally, there are works that utilize application-level information to manage data placement at the object level [6, 4, 25, 37, 44, 55, 56]. However, although these works strive to optimize the performance of memory tiering systems, their performance is limited by the inherent inaccuracy of software profiling and the high overhead of data migration using the CPU, as described in Section II. Our HeteroMem design employs a hardware-based high-resolution profiling scheme for hot and cold page detection. Meanwhile, HeteroMem uses hardware-based data movement, which achieves high bandwidth and low overhead without wasting CPU cycles and CXL link bandwidth.
VII-B Hardware-based memory tiering system
@@
There are also many works that try to optimize memory tiering systems using hardware-based methods. They manage fast memory as a memory cache [1] or use fast memory in flat memory mode [12, 47, 31, 17, 42, 35]. Intel Optane persistent memory provides memory mode [1], allowing users to use DRAM as a cache for non-volatile memory. CAMEO [12] combines memory cache mode and flat address mode. It manages fast memory in a cache-like manner, constrains data movement within associated sets, and moves data at cacheline granularity. MemPod [42] groups HBM and DRAM connected to CPU into clusters called Pods and manages data placement within each Pod. However, these hardware-based methods, proposed before the CXL era, involve intrusive modifications to CPU, making them inflexible and unable to adapt to dynamic and diverse CXL-based device-side heterogeneous memory systems. HeteroMem, on the other hand, places all its hardware logic on the device side and can be runtime configured to adapt to various configurations, offering a one-for-all solution.
@@
VIII Discussion & Future Work
Support for CXL memory pooling and sharing: Memory pooling and sharing are promising use cases of CXL [22], allowing memory to be shared between processors [13]. Memory in a memory pool can be dynamically allocated and deallocated, leading to more efficient use of memory resources. Memory sharing, on the other hand, enables each processor to read from or write to the shared memory, facilitating easy data sharing and communication between processors. The data migration process of HeteroMem is completely transparent to CPU. As a result, HeteroMem serves as a plug-in solution for both memory pooling and sharing scenarios.
Hardware overhead and Scalability: The hardware overhead of HeteroMem remains stable as the CXL-extended memory system scales. Firstly, HeteroMem’s hardware overhead increases minimally with the number of CXL devices. HeteroMem can be integrated into a CXL switch, and the number of CXL switches is significantly smaller than the number of CXL devices. Furthermore, not every CXL switch needs to be equipped with HeteroMem hardware. HeteroMem can manage the entire memory space it can access, so it only needs to be implemented in the CXL switch connected to the host CPU. Secondly, as the memory space scales, the increase in HeteroMem’s hardware overhead remains minimal. The hot page profiling component of HeteroMem utilizes Count-Min Sketch to detect hot pages, with an error bound related to the memory access sequence length rather than the memory space size. The cold page profiling component uses two bits of SRAM for every page. This structure only needs to be allocated for fast memory regions with demotion targets, which typically have a small capacity.
@@ IX Conclusion
Acknowledgments
We thank all the reviewers for their valuable comments. This work is supported by National Key Research and Development Program of China (Grant No.2023YFB4502702). This work is also supported by Beijing Natural Science Foundation (Grant No. L243001), National Natural Science Foundation of China (Grant No. 62032001), 111 Project (B18001) and the Natural Science Foundation of China (Grant No. 62332021 and 62472007).